# Static quantification of player value for fantasy basketball

### v4

@@ 1 Introduction
1.1 What is fantasy basketball?
Fantasy basketball is a game in which participants manage virtual basketball teams based on the performances of real players. According to numbers from the Fantasy Sports & Gaming Association, approximately 20 million Americans and Canadians played fantasy basketball in 2022 [1]
1.1.1 Player selection
In fantasy basketball, “managers” draft actual players to their “teams” before the NBA or WNBA season. Like in a real sports draft, no player can be selected more than once, and once a player gets picked by a manager they are removed from the draft board. Usually, the picks are made in a snaking order as depicted in Table 1. Alternatively, players can be chosen via a mock auction for which all managers start with the same amount of virtual money [2].
@@ 1.1.1 Player selection
1.1.2 Scoring
Fantasy basketball has several common scoring formats. As described by ESPN, the category-based formats are
  • Rotisserie: Rotisserie, or “Roto”, is the most common way to play fantasy basketball. In this scoring type, teams are ranked from first to last in each statistical category. Points are then awarded according to the order in each category and totaled to determine an overall score and league rank.
  • Head-to-Head: Each Category: H2H Each Category is the most common type of head-to-head play in fantasy basketball. For each scoring period, team totals are accumulated and a win, loss or tie is credited in each category based on the matchup results (i.e. 6-3-1 in a 10 category league) .
  • Head-to-Head: Most Categories: For each scoring period, team totals are accumulated in each of the categories. At the end of the scoring period, the winner is determined by which team wins the most number of categories. The end result is a win (1-0-0) , loss (0-1-0) or tie (0-0-1) . These results correspond directly to each team’s overall record [3].
@@ 1.1.2 Scoring
Table 4: Scoring with four teams, nine categories, and three weeks. Teams face each other in a round robin order
1.1.3 Position rules
@@
1.2 Static value assessment
The fantasy basketball problem is often reduced to a matter of static player values independent of context. Many fantasy basketball experts e.g. Basketball Monster publish general value estimates before each season, based on how they expect each player to perform [4]. Managers can then use these to decide which players to pick during their snake drafts. Managers can also use static value estimates for auctions, converting them to dollar values by subtracting out replacement value (the value of a player that can be added for free after the auction) then re-scaling to the amount of money available [5].
Quantifying value for category leagues is challenging because there is no obvious way to combine performances across different categories. The typical method used by Basketball Monster and others is to calculate Z-scores of expected performances by subtracting the mean for the category and dividing by standard deviation, then summing the categorical Z-scores for each player [6].
As will be shown for a simplified model of fantasy basketball, this “Z-score” method can be justified, but only as a special case of the more generally applicable “G-score” method. For the model of fantasy basketball that will be presented, G-scores are a more complete solution
@@ 2.2 Solution
It is posited that under the conditions of the simplified model of fantasy basketball laid out in Section 2.1, a novel system dubbed G-score is appropriate for quantifying player value. G-scores per category are defined by Figure 5(c), and their sum is the aggregate G-score for a player. Essentially, G-scores are the same as Z-scores except the denominators are modified to add period-to-period variance to the player-to-player-variance already included in Z-scores. Z-scores are equivalent to G-scores with $|W|=1$, removing all period-to-period variance.
@@
A detailed proof for why the G-score solves the problem is included in Appendix A. More briefly, all players contribute period-to-period variance $\tau_{M}^{2}$, and all players except the player in question contribute player-to-player variance $\sigma_{M}^{2}$. Invoking the central limit theorem, the point differential between two teams for a counting statistic can then be written as
@@
If this value is below zero, than team $A$ scores more points than team $B$ and therefore wins the category. The probability that this value is below zero is the cumulative distribution function of the normal distribution at zero. Therefore, the probability of victory is the cumulative distribution function of the distribution at zero, which is approximately
@@
This includes the formula for G-score. Applying the same procedure to the percentage statistics is more complicated, but has an equivalent result. The manager’s objective function $V$ is the sum of winning probabilities across categories, so it is
@@ 2.3 Simulation
Simulations were run to test the efficacy of G-scoring. The simulated seasons were twelve-team, thirteen-player head-to-head competitions. Each player’s performance in each simulated season was generated by randomly sampling twenty weeks from their actual performances, excluding weeks for which they were injured. Players were only included if they had at least 10 non-injured weeks. Teams were paired against each other via round robin format, as they are in real head to head competition, and winners were decided by which team had the most points by the end of the twenty week season. All games played by all players counted. Managers had access to weekly performance numbers for each player, allowing them to calculate all relevant metrics for all players. Z-scores based on the full league were used to choose $Q$, based on which managers calculated Z-scores and G-scores. The $\kappa$ factor was approximated at $1$. For the Z vs G matchup, Z-scoring was was tried at each “seat” (position in the draft order) of a snake draft, against a field of drafters using G-score. One thousand of the twenty-week simulated seasons were run for every draft seat, allowing for robust estimates of how well that strategy really would have performed in that situation with error bars no greater than
@@ 3.1 Model assumptions
Of course, the model presented in Section 2.1 to justify G-scores is not a perfect replica of how fantasy basketball works in practice. However, it does capture the main considerations, and is more feasible to solve than alternatives. This justifies it as a model worthy of study
3.1.1 Examining model assumptions
While no manager would pick players truly at random, if one picked players purely based on a ranking list with no consideration for team fit, their per-category totals would be functionally mostly random. This makes fully random selection a decent proxy for other drafters picking via ranking list, which would be much more complicated to analyze as discussed in LABEL:Comp.
The set of possible performances $W$ represents a manager’s belief about the probability distributions underlying future performance. $|W|$ can be infinite, making the probability distribution continuous, so any belief about player performance can be encoded with this model. The only sense in which this distorts real fantasy basketball is that it does not account for performance distribution changes throughout a season. See Appendix C.1 and C.2 for discussion on how to handle that complexity.
Ignoring position requirements is also not a terrible assumption for fantasy basketball, since position requirements are usually loose. See Appendix B for more discussion.
Managers generally want to maximize the expected value of their number of categories won. That is somewhat skewed from the real objective for Most Categories, but with a static ranking list there is no way of incorporating which categories a team is already weak or strong in, so maximizing for expected value is a reasonable proxy.
Finally, the idea of a value system is to use it to determine whether a player or group of players has more value than another. Ideally this calculation would consider player values in the context of each other, but that is impossible with a static system. Therefore, the only sensible way to aggregate value is to add marginal impacts of individual players
3.1.2 Computational complexity of game theory solution
Theoretically, fantasy basketball can be modeled as a game wherein managers choose their own ranking lists and draft players accordingly. That game mimics real fantasy basketball better than the model laid out in Section 2.1, since it obviates the need to assume that other players are chosen randomly. It is also a finite game, so by Nash’s theorem it has a Nash Equilibrium (an equilibrium from which no manager has incentive to deviate). Calculating this Nash Equilibrium would be a more precise way to solve the static value quantification problem. However, performing the calculation would be computationally intractable.
To calculate the Nash equilibrium, it would be necessary to know how each manager would fare under each possible strategy profile. Say that there were twelve teams with thirteen picks each, and there were $500$ total candidates. The number of possible rankings of relevant players would be
@@ 3.2 Consequences for Head-to-head and Rotisserie
Z-scores are not as problematic for Rotisserie, because weekly variations do not matter across a full season. The result is that $\tau_{M}^{2}$ is likely small relative to $\sigma_{M}^{2}$, so dropping it as Z-scores do is not as problematic. Still, using Z-scores for Rotisserie is not precisely optimal, since they ignore that some categories are harder to forecast for a full season than others
3.3 Intuition
Implicitly, Z-scores are accounting for the fact that other players are unknowns, and therefore they introduce variance in performance means. Z-scores downweight the categories that have the most uncertainty across players, and upweight the categories that are most consistent.
This logic is reasonable but also incomplete, in the sense that if player selections are unknowns, then player performances should be unknowns as well. G-scores are a version of Z-scores fixed to account for this additional uncertainty. They downweight categories that are volatile in terms of performances in the same way that Z-scores downweight categories that are volatile across players
For intuition on why this re-weighting makes sense, consider a category which has significant uncertainty across players. A Z-score advantage in that category is worth relatively little, because random chance in other player choices has a relatively large effect. There is more marginal benefit to investing in a more stable category, for which a slight advantage can translate into consistent wins. G-scoring accounts for this and makes it so that an additional point of value is worth the same for each category. When all managers are using it, it creates a sort of equilibrium, where no manager can gain an advantage by targeting less volatile categories
3.4 General utility of static ranking lists
While convenient, static ranking lists are also fundamentally suboptimal, because they preclude the possibility of adapting based on draft circumstances.
One common adaptive strategy is called punting. When a manager punts, they strategically sacrifice one category in order to gain a consistent edge in others. To punt effectively, a manager needs to adapt to the players that happen to be available and form a punting strategy around them. For example, if a player like Giannis Antetokounmpo who is generally strong but weak in free throw percentage is available, it suggests that punting free throws could be a profitable strategy. The same strategy would not work nearly as well if a player like Giannis was not available. It cannot be known beforehand whether Giannis will be available to draft, making it impossible to craft an optimal punting strategy before the draft has begun. Static ranking lists therefore cannot punt optimally.
When interpreting these results, it should be remembered both that the simulated model is incomplete, and that static ranking lists are fundamentally suboptimal. In a real draft, with managers making more optimal decisions, these results will not necessarily be applicable. Still, these simulations provide some reassurance that the basis of G-scores is sound
3.5 Simulation results
Table 6 demonstrates that drafters selecting players via Z-score consistently performed poorly against opponents using G-score, always scoring far below the baseline expected rate of $\frac{1}{12}$ or 8.33%. In the reverse scenario, G-score drafters performed much better. They won $21.4\%$ of simulated Each Category seasons, and $32.5\%$ of Most Categories seasons.
When interpreting these results, it should be remembered both that the simulated model is incomplete, and that static ranking lists are fundamentally suboptimal. In a real draft, with managers making more optimal decisions, these results will not necessarily be applicable
@@ 3.6.1 Defining Q
For simplicity, it is also possible to estimate $Q$ using a popular ranking metric
@@ 4 Conclusion
There are many areas for future work, including optimizing picks dynamically based on drafting context. Static ranking lists are fundamentally suboptimal relative to dynamic strategies so a truly sophisticated treatment of fantasy basketball would model some level of adaptation.
@@
Appendix A Detailed justification of G-score
A.1 Mathematical assumptions
@@
  • Category totals for a team can be estimated as Normal distributions due to the central limit theorem
@@
  • $\frac{X}{Y}$ can be approximated as a Normal distribution with $E\left(\frac{X}{Y}\right)\approx\frac{E(X)}{E(Y)}$. This assumes that $X$ and $Y$ are approximately Normal. independent, positive, and have low coefficient of variations [7]
@@ A.3.1 Mean of the differential
A.3.2 Variance of the differential
The variance requires the Taylor expansion, which is [8]
@@ A.3.3 Differential distribution
A.4 Total across categories
@@
Appendix B Position requirement
Center Point Guard Shooting Guard Power Forward Small Forward
Rank
1-52 13 12 10 9 8
53-104 10 11 13 11 6
105-156 7 10 11 15 8
Overall 30 34 34 36 22
Table 8: Top 156 players by G-score in 2023
If a league has particularly tight position requirements, then a value-over-replacement (VOR) approach across positions may be warranted. VOR adjusts values by estimating how valuable the best player available after the draft would be for each position, then subtracts those values from all players of those positions. It is commonly used for fantasy football, which has tight position requirement [9]
Appendix C Change in performance distributions over time
Performance distributions can change throughout a season. In lieu of a comprehensive model to account for any kind of performance drift rigorously, two simplified models are presented here to account for two major kinds of change over time- injury and random performance drift
C.1 Injuries
Sometimes, players are too injured to play for extended periods of time. As unfortunate as this is, it is a part of fantasy basketball.
The obvious way to model injury risk is to include the probability of injury into the probability distributions of player performances represented by $W$. This is not entirely satisfactory because leagues often have designed spots for injured players, allowing managers to replace them with replacement-level alternatives under certain conditions. An ideal treatment of injury risk would account for that, and ameliorate the expected effect of injury accordingly. Additionally, if a manager is concerned more with winning a league than with placing well in general, they may want to assume good health of their players, because their chances of winning would be minimal otherwise.
These complications motivate a treatment of injury risk which is parameterized and adjustable based on circumstances. One solution is scaling injury risk down by a parameter, then calculating scores with replacement value added in and scaled down by a second parameter.
In math, first player statistics are modified
$I^{\prime}=I*\upsilon$
$\mu_{c}^{\prime}=\mu_{c}*(1-I^{\prime})$
Where $\upsilon$ is a factor between zero and one, $I^{\prime}$ is an adjusted injury rate $I$, and $\mu_{c}^{\prime}$ is an adjusted mean of a counting statistic $\mu_{c}$. In this context volume statistics such as free throws attempted are included as counting statistics, since they need to be adjusted as well. Note that this adjustment should not be included in the calculation of coefficients for scoring, since it does not affect aggregate injury expectation.
Then, after scores are calculated, they are adjusted based on replacement value
$X^{\prime}=X+I^{\prime}*\psi*R$
Where $\psi$ is the expected fraction of games for which an injured player can be replaced, $R$ is replacement value, and $X^{\prime}$ is a modified score $X$.
Replacement-level value can be approximated in several ways. One is to spread value evenly, except with a negative for turnovers. So for example if the replacement value is $-3$, by category it is broken down by $R=-\frac{3}{7}$ for non-turnover categories and $R=\frac{3}{7}$ for turnovers. The sum is $-3$ as needed
C.2 Performance Drift
Player performances inevitably drift over time. When attentive managers are allowed to replace under-performing players, they can replace low-level players that dip below replacement-level value. This dilutes the value of low-level players who are likely to be replaced, in a way that can be analyzed mathematically
C.2.1 A simple model of performance drift
Assume that the replacement value remains the same over the course of a season. This is intuitively reasonable; individual players may go up or down but the value of the player at rank e.g. 157 should be relatively consistent.
Further, assume that each player’s value is perturbed by noise $\epsilon$ with mean zero and standard deviation $\sigma$ by some point during the season. If $\epsilon$ drops the player’s value below the replacement value, then they are not worth having on a team and will be dropped in favor of a player with exactly the replacement value.
Consider the expected performance value of a player above replacement to be $\mu$, with perturbation $\epsilon_{f}$. Then, their perturbed value $F$ is
$F=\begin{cases}\mu+\epsilon_{f},&\text{if }\mu+\epsilon_{f}\geq 0\\ 0,&\text{otherwise}\end{cases}$
Without any investment into a particular roster slot, a manager can always choose a replacement-level player for essentially no cost. Players with exactly replacement value before perturbation ($\mu=0$) have post-perturbation value
$S=\begin{cases}\epsilon_{s},&\text{if }\epsilon_{s}\geq 0\\ 0,&\text{otherwise}\end{cases}$
Since picking a player removes a spot that could be used for a costless replacement-level player or “streamer”, the real value of picking a player is $E[F-S]=E[F]-E[S]$
C.2.2 SAVOR
It is posited that a novel metric called SAVOR is equivalent to $E[F]-E[S]$ given the assumptions. SAVOR-adjusted player value given starting point $\mu$ is
$\mu*\phi(\frac{\mu}{\sigma})-\frac{\sigma}{\sqrt{2\pi}}\left(1-e^{\frac{-\mu^{2}}{2\sigma^{2}}}\right)$
Where $\phi$ is the cumulative distribution function of a standard Normal.
This adjustment is monotonic with respect to $\mu$, so it only has relevance to auctions, where value is important in addition to order. Table 9 shows an example of its application. As expected, the adjustment dilutes the value of low-level players, to a degree that increases with $\sigma$
Pre-Savor Post-SAVOR
$\mu$ $\sigma=0$ $\sigma=0.1$ $\sigma=1$
Value Percent Value Percent Value Percent Value Percent
0 0% 0 0% 0 0% 0 0%
1 10% 1 10% 0.96 9.76% 0.68 8.05%
2 20% 2 20% 1.96 19.92% 1.61 18.94%
3 30% 3 30% 2.96 30.08% 2.60 30.62%
4 40% 4 40% 3.96 40.24% 3.60 42.38%
Table 9: Results of a SAVOR calculation, re-normalized to sum to 100
C.2.3 Justifying SAVOR
With $\epsilon_{f}$ and $\epsilon_{s}$ defined as normal errors with standard deviation $\sigma$,
$F=\begin{cases}\mu+\epsilon_{f},&\text{if }\mu+\epsilon_{f}\geq 0\\ 0,&\text{otherwise}\end{cases}$
$S=\begin{cases}\epsilon_{s},&\text{if }\epsilon_{s}\geq 0\\ 0,&\text{otherwise}\end{cases}$
Using the definition of the expected value and the probability density of the normal distribution,
[…]
Changing the variable in the expression for $E[F]$ from $x$ to $y+\mu$ yields
[…]
Therefore
[…]
The second part of the expression can be identified as $\mu*\phi(\frac{\mu}{\sigma})$. This simplifies to
[…]
Substituting $-\sigma^{2}*d(-\frac{y^{2}}{2\sigma^{2}})$ for $ydy$ yields
$E[F]-E[S]=\mu*\phi(\frac{\mu}{\sigma})-\int_{-\mu}^{0}\frac{\sigma}{\sqrt{2\pi}}e^{\frac{-y^{2}}{2\sigma^{2}}}d(-\frac{y^{2}}{2\sigma^{2}})$
Upon integration, this yields
$E[F]-E[S]=\mu*\phi(\frac{\mu}{\sigma})-\frac{\sigma}{\sqrt{2\pi}}\left(e^{\frac{-y^{2}}{2\sigma^{2}}}|_{-\mu}^{0}\right)=\mu*\phi(\frac{\mu}{\sigma})-\frac{\sigma}{\sqrt{2\pi}}\left(1-e^{\frac{-\mu^{2}}{2\sigma^{2}}}\right)$

### v5

@@ 1 Introduction
Fantasy basketball is a game in which participants manage virtual basketball teams linked to the performances of real players. According to numbers from the Fantasy Sports & Gaming Association, approximately 20 million Americans and Canadians played fantasy basketball in 2022 (FSGA, 2023).
Most fantasy basketball leagues have category-based scoring (Bartilotta, 2023). Despite the broad popularity of fantasy basketball, the literature on so-called category leagues is scant. Quantitatively, the fantasy basketball community relies on a heuristic metric called Z-score to estimate the values of different players for them (Kelly, 2021). As will be shown, the metric can be justified under certain assumptions, but is missing a crucial element
1.1 Player selection
In fantasy basketball, “managers” draft actual players to their “teams” before the NBA or WNBA season. Like in a real sports draft, no player can be selected more than once, and once a player gets picked by a manager they are removed from the draft board. Usually, the picks are made in a snaking order as depicted in Table 1. Alternatively, players can be chosen via a mock auction for which all managers start with the same amount of virtual money (ESPN, 2023).
@@ 1.1 Player selection
1.2 Scoring
Most fantasy basketball leagues are category leagues, meaning that they score managers based on how many categories they win against each other (Bartilotta, 2023). The three main formats for category leagues are
  • Rotisserie: Teams are ranked according to aggregate performance in each statistical category over the course of a season. The team with the best average ranking across categories wins
  • Head-to-Head: Each Category: Teams are matched up against opponents each week and their category-level results are tabulated. E.g. if a team wins six categories, loses two, and ties one, they score 6-2-1. Teams are ranked at the end of the season by aggregate records
  • Head-to-Head: Most Categories: Teams are matched up against opponents each week and the team that wins a majority of categories wins the week. Teams are ranked at the end of the season by number of weeks won (ESPN, 2023)
@@ 1.2 Scoring
Table 4: Scoring with four teams, nine categories, and three weeks
1.3 Position rules
@@
2 Z-scores for static value
A core component of fantasy basketball strategy is having an informed opinion of which players are best to draft.
To this end, many fantasy basketball experts e.g. Basketball Monster publish estimates of player value before each season, based on how they expect players to perform (RotoMonster, no date). These value estimates do not depend on draft circumstances, which make them convenient and easy to use. Managers can apply them during snake drafts by taking players with the highest available values, without needing to adjust for previously drafted players. Managers can also use them for auctions, converting them to dollar values by subtracting out replacement value (the value of a player that can be added for free after the auction) then re-scaling to the amount of money available (Roberts, 2014). The resulting dollar values can be used as benchmarks for how much they should be willing to spend on players.
Naturally, experts who make value estimates often try to do so in a mathematical way. This is non-trivial for category leagues, because there is no obvious way to aggregate expected performances across categories.
Z-scoring is the traditional method for static value quantification used by Basketball Monster and others (Kelly, 2021). It is a three-step process
  1. 1.
    For percentage statistics, multiply expected performances by expected volume, divided by average volume
  2. 2.
    For all categories, subtract out the mean and divide by the standard deviation
  3. 3.
    Sum up all of the category-level scores to get scores for each player
This approach is attractive because it ensures that the distributions per category all have a mean of zero and standard deviation of one, which intuitively seems reasonable for a scoring system. However, being intuitively reasonable is not proof that Z-scores are optimal. Understanding the sense in which Z-scores are or are not optimal requires mathematical analysis
@@ 3 Justifying G-scores
Under a synthetic version of the static value problem which is described in this section,, Z-scores are close to optimal but not fully general. A novel metric dubbed G-score fixes Z-score by extending it to incorporate uncertainty in player performances
@@ 3.2 Solution
It is posited that G-scores, as defined by Figure 5(c), are optimal under the conditions of the simplified model of fantasy basketball laid out in Section 3.1. Essentially, G-scores are the same as Z-scores except the denominators are modified to add period-to-period variance to the player-to-player-variance already included in Z-scores. Z-scores are equivalent to G-scores with $|W|=1$, removing all period-to-period variance
@@
3.3 Proof
A detailed proof for why the G-score solves the problem of Section 3.1 is included in the appendix. More briefly, all players contribute period-to-period variance $\tau_{M}^{2}$, and all players except the player in question contribute player-to-player variance $\sigma_{M}^{2}$. Invoking the central limit theorem, the point differential between two teams for a counting statistic can then be written as
@@ 3.3 Proof
If this value is below zero, than team $A$ scores more points than team $B$ and therefore wins the category. The probability that this value is below zero is its cumulative distribution function at zero. Invoking the Taylor Series approximation for the CDF of the normal distribution and some algebraic manipulation, that probability is
@@
This includes the formula for G-score for counting statistics. Applying the same procedure to the percentage statistics is more complicated, but results in the equivalent
$\frac{1}{2}\left[1+\frac{1}{\sqrt{\pi\left(N-\frac{1}{2}\right)}}*\frac{\frac{\mu_{A}\left(p\right)}{\mu_{A}}\left(\mu_{R}\left(p\right)-\mu_{R}\right)}{\sqrt{\sigma_{R}^{2}+\frac{2N}{2N-1}\tau_{R}^{2}}}\right]$
Which includes the G-score for percentage statistics. The manager’s objective function $V$ is the sum of winning probabilities across categories, so it is
@@ 3.4 Simulation
Simulations were run to test the efficacy of G-scoring. The simulated seasons were twelve-team, thirteen-player head-to-head competitions. Each player’s performance in each simulated season was generated by randomly sampling twenty weeks from their actual performances, excluding weeks for which they were injured. Players were only included if they had at least 10 non-injured weeks. Teams were paired against each other and winners were decided by which team had the most points by the end of the twenty week season. All games played by all players counted. Managers had access to weekly performance numbers for each player, allowing them to calculate all relevant metrics for all players. Z-scores based on the full league were used to choose $Q$, based on which managers calculated Z-scores and G-scores. The $\kappa$ factor was approximated at $1$. For the Z vs G matchup, Z-scoring was was tried at each “seat” (position in the draft order) of a snake draft, against a field of drafters using G-score. One thousand of the twenty-week simulated seasons were run for every draft seat, allowing for robust estimates of how well that strategy really would have performed in that situation with error bars no greater than
@@
The results are shown in Table 6. It demonstrates that drafters selecting players via Z-score consistently performed poorly against opponents using G-score, always scoring far below the baseline expected rate of $\frac{1}{12}$ or 8.33%. In the reverse scenario, G-score performed much better. It won $21.4\%$ of simulated Each Category seasons, and $32.5\%$ of Most Categories seasons.
@@ 4.1 Model assumptions
Of course, the model presented in Section 3.1 to justify G-scores is not a perfect replica of how fantasy basketball works in practice. The degree to which the model represents true fantasy basketball merits discussion
4.1.1 Players chosen randomly
While no manager would pick players truly at random, if all managers picked players purely based on a single ranking list with no consideration for team fit, it stands to reason that their per-category totals would be close to random. A manager’s category totals would result from the characteristics of the players that happened to be at the top of the ranking list for their draft slots.
The simulation results provide evidence that the assumption of players being chosen randomly was not problematic for G-scores, relative to the alternative of managers choosing players via ranking lists. They had managers choosing players based on ranking lists rather than randomly, and the G-score metric still worked well.
Theoretically, one could explicitly model fantasy basketball as a game wherein managers choose their own ranking lists and draft players accordingly. It would be possible to determine exactly which players each manager would end up with, because the draft process would be entirely deterministic. This conception of the game obviates the need to assume that other players are chosen randomly. It is also a finite game, so by Nash’s theorem it has a Nash Equilibrium (an equilibrium from which no manager has incentive to deviate). However, calculating the equilibrium would be computationally intractable. To calculate the Nash equilibrium, it would be necessary to know how each manager would fare under each possible strategy profile. Say that there were twelve teams with thirteen picks each, and there were $500$ total candidates. The number of possible rankings of relevant players would be
@@ 4.1.1 Players chosen randomly
4.1.2 Performances sampled from distribution
The set of possible performances $W$ represents a manager’s belief about the probability distributions underlying future performance. $|W|$ can be infinite, making the probability distribution continuous, so any belief about player performance can be encoded with this model.
The only sense in which this distorts real fantasy basketball is that it does not account for players having their performance distributions change throughout a season. These changes can be significant, particularly if they are due to injury. Future work could consider their effects on value quantification
4.1.3 No position requirement
Center Point Guard Shooting Guard Power Forward Small Forward
Rank
1-52 13 12 10 9 8
53-104 10 11 13 11 6
105-156 7 10 11 15 8
Overall 30 34 34 36 22
Table 7: Top 156 players by G-score in 2023
If a league has particularly tight position requirements, then a value-over-replacement (VOR) approach across positions may be warranted. VOR adjusts values by estimating how valuable the best player available after the draft would be for each position, then subtracts those values from all players of those positions. It is commonly used for fantasy football, which has tight position requirement (Fantasy Pros., 2017)
4.1.4 Objective is number of categories won
The expected value of number of categories won is not a perfect proxy for what managers want.
For one thing, it is obvious that this objective function is not correct for Most Categories. In that format, a manager would rather have a $100\%$ chance of winning five categories than a $99\%$ chance of winning all of them, even if they would win fewer categories on average.
Another issue is that managers generally want to win their leagues or get close, which is not equivalent to maximizing the expected value of number of categories won. A manager with this preference would be inclined towards risky strategies that increase the probability of a strong performance while being sub-optimal for expected value.
Neither of these issues is straightforward to incorporate into a value quantification model. With a static ranking list, there is no way of considering which categories a team is already weak or strong in, making it awkward to strategize around winning a majority of categories. And calculating the probability of winning a league would be much more complicated than calculating the probability of winning a single match-up.
Expected value of number of categories won is a reasonable compromise, because it gets at the main thrust of a managers’ goal while facilitating a straightforward approach to modeling
4.1.5 Player value is marginal relative to objective
The idea of a value system is to use it to determine whether a player or group of players has more value than another. Ideally this calculation would consider player values in the context of each other, but that is impossible with a static system. Therefore, the only sensible way to aggregate value is to add marginal impacts of individual players
4.2 General utility of static ranking lists
Granting that the model described in Section 4.1 represents the static value problem fairly, it should still be kept in mind that choosing players via a static ranking list is fundamentally sub-optimal. Drafting that way precludes the possibility of adapting based on draft circumstances, which is often advantageous.
One common adaptive strategy is called punting. When a manager punts, they strategically sacrifice one category in order to gain a consistent edge in others. To punt effectively, a manager needs to adapt to the players that happen to be available and form a punting strategy around them (Mamone, 2022). For example, if a player like Giannis Antetokounmpo who is generally strong but weak in free throw percentage is available, it suggests that punting free throws could be a profitable strategy. The same strategy would not work nearly as well if a player like Giannis was not available. It cannot be known beforehand whether Giannis will be available to draft, making it impossible to craft an optimal punting strategy before the draft has begun. Static ranking lists therefore cannot punt optimally.
Another limitation of static ranking lists is that they can lead to unbalanced teams. For example, if the strongest players available to a drafter always happen to have extremely high rebound numbers, the drafter using a static ranking list would take all of them. This might allow them to dominate in rebounds but would not be optimal overall, since the team would be weak in other categories. Winning one category by a large margin and losing all others would result in a poor overall performance.
This is not to say that static value quantification is inherently useless. It can be useful for a context wherein managers are not capable of applying advanced strategies. Also, it can serve as a starting point for managers formulating more advanced strategies
@@ 4.3 Consequences for head-to-head and Rotisserie
Intuitively, this makes sense because volatile categories like steals are hard to win even with investment. Under the Z-score framework, an extra point of steals is not worth as much as an extra point of assists, because the value of the investment in steals is diluted by natural variance. G-scores account for this and make it so that scores for all categories have equivalent value.
Z-scores are not as problematic for Rotisserie, because weekly variations wash out to a degree across a full season. Still, using Z-scores for Rotisserie is not precisely optimal, since they ignore that some categories are harder to forecast for a full season than others
@@ 4.4.1 Defining Q
It is also possible to estimate $Q$ using a popular ranking metric. This might be preferable in practice because G-scores are not necessarily reflective of how other managers actually value players
@@ 5 Conclusion
There are many areas for future work, including optimizing picks dynamically based on drafting context. Static ranking lists are fundamentally sub-optimal relative to dynamic strategies so a truly sophisticated treatment of fantasy basketball would model some level of adaptation.
@@
Appendix- Detailed justification of G-score
Mathematical assumptions
@@
  • Category totals for a team can be estimated as Normal distributions due to the central limit theorem. Since there are usually $\approx 26$ relevant players, whose performances are within some standard range (fluctuations from 20 to 40 points per game are normal, but nobody is scoring 100), this is intuitively reasonable
@@
  • $\frac{X}{Y}$ can be approximated as a Normal distribution with $E\left(\frac{X}{Y}\right)\approx\frac{E(X)}{E(Y)}$. This assumes that $X$ and $Y$ are approximately Normal. independent, positive, and have low coefficient of variations (Diaz-Frances, 2012)
@@
Variance of the differential
The variance requires the Taylor expansion, which is (Benaroya, 2005)
@@
Total across categories